# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

### v2

@@ 5. Homology Theories on Real $C^{*}$-algebras
    $0\longrightarrow{\mathcal{K}}_{R}\longrightarrow\mathcal{T}_{{\mathbb{R}}0}\longrightarrow V{\mathbb{R}}\longrightarrow 0.$
@@
    $0\longrightarrow A\otimes{\mathcal{K}}_{R}\longrightarrow\mathcal{T}_{{\mathbb{R}}0}\otimes A\longrightarrow VA\longrightarrow 0.$
@@
and this implies [^24] This is a matter of pure algebra: suppose given an exact sequence of abelian groups $G_{1}\overset{\alpha_{1}}{\longrightarrow}G_{2}\overset{\alpha_{2}}{\longrightarrow}G_{3}\overset{\alpha_{3}}{\longrightarrow}G_{4}\overset{\alpha_{4}}{\longrightarrow}G_{5}$ with $\alpha_{1}$ and $\alpha_{4}$ isomorphisms. Then $Ker(\alpha_{2})\cong Im(\alpha_{1})=G_{2}$ since $\alpha_{1}$ is onto, which says that $\alpha_{2}\equiv 0$. Similarly, $Im(\alpha_{3})=Ker(\alpha_{4})=0$ since $\alpha_{4}$ is 1-1, which says that $\alpha_{3}\equiv 0$. Then $\alpha_{2}=\alpha_{3}\equiv 0$ implies that $G_{3}=0$. that $h_{n}(\mathcal{T}_{{\mathbb{R}}0}\otimes A)=0$ for all $n$, establishing Part (3).
@@
The following corollary describes both suspension (which also is available in the complex $K$-theory situation) and desuspension (which is totally without precedent in the complex $K$-theory world).
@@
For any homology theory $h_{*}$ on real $C^{*}$-algebras and for any real $C^{*}$-algebra $A$, there is a natural suspension isomorphism
@@
and a natural desuspension isomorphism
@@
Next we turn our attention to the creation of new homology theories. This technique has been used before, for example in [75] to define $K_{*}(A;{\mathbb{Z}}_{n})$, complex $K$-theory with coefficients ${\mathbb{Z}}_{n}$ from $K_{*}(-)$.
@@
Why discuss cofibre homology theories if they seem unrelated to $KO$-theory? The basic point to understand is that, in classical algebraic topology, most homology and cohomology theories are in fact only cofibre theories. Topological $K$ and $KO$-theories are actually quite exceptional. We anticipate that in years to come that there will arise other theories to help us understand real $C^{*}$-algebras. Given the precedent in classical algebraic topology, there is no reason to suppose that these theories will satisfy the axioms that we have stated for homology theories, but they will be cofibre theories. It is already the case that cofibrations252525known there as ”Schochet fibrations” as discussed in Section 4. have appeared in several papers, for instance in Bunke [31].
@@ 6. $KO_{*}$-theory for real $C^{*}$-Algebras: Definitions and Basic Properties
Let $A$ be a unital ring (for example, a unital real $C^{*}$-algebra). A finitely-generated free $A$-module is a module of the form $A\oplus\dots\oplus A$, with a finite number (say $k$) of copies of $A$. Think of this as the identity function
@@
but in general the map will not be 1-1 or onto. Let us denote by $P$ the image of the projection. Then $P$ is a finitely-generated $A$-module. The orthogonal complement of $p$ (call it $q$) has image that is orthogonal to $P$ and $P\oplus Q\cong A^{k}$. This is exactly what is required to say that $P$ is a finitely-generated projective $A$-module. Again, the process may be reversed.
@@
Given any homomorphism $A\to B$, there is induced a homomorphism
@@
Thus $KO_{0}$ is a covariant functor from unital $C^{*}$-algebras and maps to abelian groups.
@@
For each real $C^{*}$-algebra $A$ there is a natural complexification map
@@
which produces a natural transformation of theories. To start with, suppose that $P$ is a finitely generated projective $A$-module with associated isomorphism
$P\oplus Q\cong A^{k}.$
By complexifying we obtain the isomorphism
$P_{\mathbb{C}}\oplus Q_{\mathbb{C}}\cong(A_{\mathbb{C}})^{k}$
and hence $P_{\mathbb{C}}$ is a finitely generated projective $A_{\mathbb{C}}$-module. This induces a natural map
@@
for $A$ unital by sending $[P]$ to $P_{\mathbb{C}}]$. The map is obviously a natural homomorphism. As the definition of the $KO$ and $K$-theory groups for nonunital algebras are essentially identical, it is easy to show that
$c:KO_{0}(A)\longrightarrow K_{0}(A_{\mathbb{C}})$
is well-defined and well-behaved for all real $C^{*}$-algebras.
We can describe this map in a different way. For any real $C^{*}$-algebra there is an inclusion map, which is a homomorphism $A\hookrightarrow A_{\mathbb{C}}$. If we regard $A_{\mathbb{C}}$ as a real $C^{*}$-algebra this induces a natural transformation
@@
and so the composition gives
@@
and we shall see that this induces a short exact sequence
@@
We note that it is straight-forward to prove313131using CW-complexes, which we will discuss in §9. that the natural maps
@@
Then each $KO_{n}(-)$ is a covariant functor from real $C^{*}$-algebras to abelian groups. We write $KO_{*}(A)$ for the graded abelian group $\{KO_{n}(A):n\geq 0\}$. The same construction for complex $C^{*}$-algebras yields $K_{j}(A)$.
For many purposes it is convenient to use the fact that $O_{n}(A)$, the group of orthogonal $n\times n$ matrices, is a strong deformation retraction of $GL_{n}(A)$. [^33] For example, the set $O_{1}({\mathbb{R}})=\{\pm 1\}$ is a strong deformation retraction of $GL_{1}({\mathbb{R}})={\mathbb{R}}-\{0\}$, and $O_{2}$ (rotations and reflections of the plane - topologically two circles) is a strong deformation retraction of $GL_{2}({\mathbb{R}})$. The compact Lie group $O_{n}({\mathbb{R}})$ is naturally contained in $GL_{n}({\mathbb{R}})$ and there is a strong deformation retraction343434also known as Gram-Schmidt orthogonalization which deforms $GL_{n}({\mathbb{R}})$ down to $O_{n}({\mathbb{R}})$. Thus the two spaces are homotopy equivalent. This implies that $O_{\infty}({\mathbb{R}})$ has the same homotopy type as $GL_{\infty}({\mathbb{R}})$, and an analogous argument demonstrates that $O_{\infty}(A)$ has the same homotopy type as $GL_{\infty}(A)$ for any real (or complex!) $C^{*}$-algebra $A$. Thus we may as well define
@@
This result has been “well-known” for fifty years or so, but it is remarkably difficult to put together a rigorous proof. [^35] Jon Rosenberg (private communication) suggested the following approach. First, define bigraded $K$-theory via Clifford algebras and use the periodicity for Clifford algebras to show that $KO_{*}(A)$ is periodic of period 8. The resulting homology theory may be represented by an infinite loop space, and you then show that this infinite loop space may be taken to be determined by $GL_{\infty}(A)$. For information on Clifford algebras and periodicity please see future sections. The technical details of this approach are well beyond the scope of this work . Schröder (cf. Schröder, [76] Theorem 1.4.6 ) briefly discusses the proof of this theorem and then refers to Karoubi’s book [52] III, 1.6 (it is actually 1.11). The book does prove periodicity for $K_{*}$ but not for $KO_{*}$. For a detailed proof of that you have to go to an earlier paper of Karoubi, namely [45].
@@
    Now $SO_{2}\cong S^{1}$ and so $\pi_{1}(SO_{2})\cong{\mathbb{Z}}$, but $SO_{3}\cong RP^{3}$ as a topological space (where $RP^{3}$ is real projective space) and hence $\pi_{1}(SO_{3})={\mathbb{Z}}_{2}$. For $n>2$, any map $S^{1}\to SO_{n}$ is homotopic373737by the Cellular Approximation Theorem, cf. G. Whitehead [86] p. 77 and see §9 ahead. to a map taking values in $SO_{3}$ and hence $\pi_{1}(SO_{n})={\mathbb{Z}}_{2}$ for $n>2$, confirming that
@@ 7. External Products, the Künneth Theorem, and Internal Products
    $\mu_{O}\colon KO_{m}(A)\otimes KO_{n}(B)\rightarrow KO_{m+n}(A\otimes B).$
@@
    $\mu_{U}\colon K_{m}(A)\otimes K_{n}(B)\rightarrow K_{m+n}(A\otimes B).$
@@
Künneth theorems deal in general with determining $h_{*}(A\otimes B)$ in terms of $h_{*}(A)$ and $h_{*}(B)$. One has to put some constraints on the algebras in order for the theorem to hold. The full Künneth Theorem for $K$-theory involves some serious homological algebra. For $K_{*}(A)$ it was established by Schochet [73] (1984) where the bootstrap category was first introduced, and for $KO_{*}(A)$ it is due to Boersema [15] (2002), utilizing united $K$-theory. The following results are special cases of these general theorems. Anticipating events, we also state the corresponding results for the $K$-theory of topological spaces. See Section 11 for $K^{*}(X)$.
@@
    Suppose that $A$ and $B$ are complex $C^{*}$-algebras, $A$ is in the bootstrap category (which includes all commutative complex $C^{*}$-algebras) and $K_{*}(B)$ is torsion-free. Then the natural external product map induces an isomorphism of graded groups [^39] If $G_{*}$ and $H_{*}$ are graded abelian groups then $(G_{*}\otimes H_{*})_{5}$ (for example) is understood to be the direct sum of the groups $G_{j}\otimes H_{5-j}$. This makes sense as long as only a finite number of these groups are non-zero. Bott periodicity makes things more complicated. So let us stipulate that whenever we are tensoring $K$-theory groups, that $\in\{0,1\}$ in the complex case and $\in\{0,1,\dots 6,7\}$ in the real case.
@@
    Suppose that $A$ and $B$ are real $C^{*}$-algebras, $A_{\mathbb{C}}$ is in the bootstrap category (this includes all commutative real $C^{*}$-algebras $A$), and $B$ has the property that $K_{*}(B_{\mathbb{C}})$ is torsion-free. Then the natural external product map induces an isomorphism of graded $KO_{*}({\mathbb{R}})$-modules [^40] We are casually assuming that $KO_{*}({\mathbb{R}})$ is a graded commutative ring that acts on $KO_{*}(A)$ for any real $C^{*}$-algebra $A$. We will establish this rigorously subsequently. Actually the analogous statement in the complex case is also true, but since $K_{*}({\mathbb{C}}))$ is just $\mathbb{Z}$} and 0 in degrees 0 and $1$ respectively, the action has little significance here.
@@
    This includes cases where $B=C^{{\mathbb{R}}}(iS^{1})$, $C_{o}^{{\mathbb{R}}}(i{\bf{R}})$, $\real(S^{n})$. In particular, if $X$ and $Y$ are compact spaces and $Y$ has the homotopy type of a sphere or a product of spheres414141or a finite or countable disjoint union of such spaces, then there is a natural isomorphism
@@
We shall describe the ring structure of $KO_{*}({\mathbb{R}})$ and $KO_{*}({\mathbb{R}})\otimes{\mathbb{Q}}$ in some detail in Section 15. Note that
$KO_{n}({\mathbb{R}})\otimes{\mathbb{Q}}=\begin{cases}{\mathbb{Q}}&n=0,\,4\\ 0&\text{otherwise}\end{cases}$
It turns out that $KO_{n}(A)\otimes{\mathbb{Q}}$ is periodic of period $4$.
@@
Since $H\otimes H\cong M_{r}({\mathbb{R}})$ as remarked in §1, this simplifies to

### v3

@@ 5. Homology Theories on Real $C^{*}$-algebras
    $0\longrightarrow{\mathcal{K}}_{\mathbb{R}}\longrightarrow\mathcal{T}_{{\mathbb{R}}0}\longrightarrow V{\mathbb{R}}\longrightarrow 0.$
@@
    $0\longrightarrow A\otimes{\mathcal{K}}_{\mathbb{R}}\longrightarrow\mathcal{T}_{{\mathbb{R}}0}\otimes A\longrightarrow VA\longrightarrow 0.$
@@
and this implies [^25] This is a matter of pure algebra: suppose given an exact sequence of abelian groups $G_{1}\overset{\alpha_{1}}{\longrightarrow}G_{2}\overset{\alpha_{2}}{\longrightarrow}G_{3}\overset{\alpha_{3}}{\longrightarrow}G_{4}\overset{\alpha_{4}}{\longrightarrow}G_{5}$ with $\alpha_{1}$ and $\alpha_{4}$ isomorphisms. Then $Ker(\alpha_{2})\cong Im(\alpha_{1})=G_{2}$ since $\alpha_{1}$ is surjective, which says that $\alpha_{2}\equiv 0$. Similarly, $Im(\alpha_{3})=Ker(\alpha_{4})=0$ since $\alpha_{4}$ is injective, which says that $\alpha_{3}\equiv 0$. Then $\alpha_{2}=\alpha_{3}\equiv 0$ implies that $G_{3}=0$. that $h_{n}(\mathcal{T}_{{\mathbb{R}}0}\otimes A)=0$ for all $n$, establishing Part (3).
@@
The following corollary describes both suspension, which also is available in the complex $K$-theory world, and desuspension, which is totally hidden in the complex $K$-theory world, since
$VA\otimes{\mathbb{C}}\cong SA\otimes{\mathbb{C}}=S(A\otimes{\mathbb{C}}),$
a fact related to Bott Periodicity.
Here’s a simple way to emphasize the relationship between suspension and desuspension.
@@
Let $h_{*}$ be a homology theory on real $C^{*}$-algebras. Then for each real $C^{*}$-algebra $A$:
  1. (1)
    There is a natural suspension isomorphism
@@
    induced from a natural short exact sequence
    $0\to SA\longrightarrow(contractible)\longrightarrow A\to 0.$
  2. (2)
    There is a natural desuspension isomorphism
@@
    induced from a natural short exact sequence
    $0\to A\otimes{\mathcal{K}}_{\mathbb{R}}\longrightarrow(contractible)\longrightarrow VA\to 0$
Next we turn our attention to the creation of new homology theories. This technique has been used before, for example in [75] to build $K_{*}(A;{\mathbb{Z}}_{n})$, complex $K$-theory with coefficients ${\mathbb{Z}}_{n}$ from $K_{*}(-)$.
@@
Why discuss cofibre homology theories if they seem unrelated to $KO$-theory? The basic point to understand is that, in classical algebraic topology, most homology and cohomology theories are in fact only cofibre theories. Topological $K$ and $KO$-theories are actually quite exceptional. We anticipate that in years to come that there will arise other theories to help us understand real $C^{*}$-algebras. Given the precedent in classical algebraic topology, there is no reason to suppose that these theories will satisfy the axioms that we have stated for homology theories, but they will be cofibre theories. It is already the case that cofibrations262626known there as ”Schochet fibrations” as discussed in Section 4. have appeared in several papers, for instance in Bunke [31].
@@ 6. $KO_{*}$-theory for Real $C^{*}$-Algebras: Definitions and Basic Properties
Let $A$ be a unital ring (for example, a unital real $C^{*}$-algebra). In the following, we choose to work with left $A$-modules throughout: one must choose left or right at the outset. A finitely-generated free $A$-module is a left $A$-module of the form $A\oplus\dots\oplus A$, with a finite number (say $k$) of copies of $A$. Think of this as the identity function
@@
but in general the map will not be injective or surjective. Let us denote by $P$ the image of the projection. Then $P$ is a finitely-generated $A$-module. The orthogonal complement of $p$ (call it $q$) has image that is orthogonal to $P$ and $P\oplus Q\cong A^{k}$. This is exactly what is required to say that $P$ is a finitely-generated projective $A$-module. Again, the process may be reversed.
@@
Suppose that $f:A\to B$ is a homomorphism of real $C^{*}$-algebras. If $p$ is a projection in $M_{n}(A)$ then $f(p)$ is a projection in $M_{n}(B)$. The map $p]\mapsto[f(p)]$ is well-defined on equivalences classes and respects the operators, so there is an induced homomorphism
@@
  3. (3)
    If $f_{0},f_{1}:A\to B$ are homotopic, then
    $(f_{0})_{*}=(f_{1})_{*}:KO_{0}(A)\longrightarrow KO_{0}(B).$
Thus $KO_{0}$ is a homotopy-invariant covariant functor from unital $C^{*}$-algebras and maps to abelian groups.
@@
For each real $C^{*}$-algebra $A$ there is an inclusion map, which is a homomorphism $A\hookrightarrow A_{\mathbb{C}}$. This induces a natural complexification map
@@
which produces a natural transformation of theories. It has an easy description. If $p$ is a projection in $A$ then it is also a projection in $A_{\mathbb{C}}$. If $p$ and $q$ are unitarily equivalent projections in $A$ then this also holds in $A_{\mathbb{C}}$. Passing to $M_{n}(A)$ and $M_{n}(A_{\mathbb{C}})$, the same holds. Therefore there is a homomorphism
@@
defined by $c([p])=[p]$, which is well-defined and well-behaved for all real $C^{*}$-algebras.
We can describe this map in a different way. If we regard $A_{\mathbb{C}}$ as a real $C^{*}$-algebra this induces a natural transformation
@@
and so the composition gives a natural homomorphism
@@
and we shall see later that this induces a short exact sequence
@@
We note that it is straight-forward to prove323232using CW-complexes, which we will discuss in §9. that the natural maps
@@
Then each $KO_{n}(-)$ is a covariant functor from real $C^{*}$-algebras to abelian groups. We write $KO_{*}(A)$ for the graded abelian group $\{KO_{n}(A):n\geq 0\}$. The same construction for complex $C^{*}$-algebras yields $K_{n}(A)$.
For many purposes it is convenient to use the fact that $O_{n}(A)$, the group of orthogonal $n\times n$ matrices, is a strong deformation retraction of $GL_{n}(A)$. [^34] For example, the set $O_{1}({\mathbb{R}})=\{\pm 1\}$ is a strong deformation retraction of $GL_{1}({\mathbb{R}})={\mathbb{R}}-\{0\}$, and $O_{2}$ (rotations and reflections of the plane - topologically two circles) is a strong deformation retraction of $GL_{2}({\mathbb{R}})$. The compact Lie group $O_{n}({\mathbb{R}})$ is naturally contained in $GL_{n}({\mathbb{R}})$ and there is a strong deformation retraction353535also known as Gram-Schmidt orthogonalization which deforms $GL_{n}({\mathbb{R}})$ down to $O_{n}({\mathbb{R}})$. Thus the two spaces are homotopy equivalent. This implies that $O_{\infty}({\mathbb{R}})$ has the same homotopy type as $GL_{\infty}({\mathbb{R}})$, and an analogous argument demonstrates that $O_{\infty}(A)$ has the same homotopy type as $GL_{\infty}(A)$ for any real (or complex!) $C^{*}$-algebra $A$. Thus we may as well define
@@
This result has been “well-known” for fifty years or so, but it is remarkably difficult to put together a rigorous proof. [^36] Jon Rosenberg (private communication) suggested the following approach. First, define bigraded $K$-theory via Clifford algebras and use the periodicity for Clifford algebras to show that $KO_{*}(A)$ is periodic of period 8. The resulting homology theory may be represented by an infinite loop space, and you then show that this infinite loop space may be taken to be determined by $GL_{\infty}(A)$. For information on Clifford algebras and periodicity please see future sections. The technical details of this approach are well beyond the scope of this monograph. Schröder (cf. Schröder, [76] Theorem 1.4.6 ) briefly discusses the proof of this theorem and then refers to Karoubi’s book [52] III, 1.6 (it is actually 1.11). The book does prove periodicity for $K_{*}$ but not for $KO_{*}$. For a detailed proof of that you have to go to an earlier paper of Karoubi, namely [45].
@@
    Alternatively, we can describe the non-trivial element of $KO_{1}({\mathbb{R}})={\mathbb{Z}}_{2}$ in terms of a projection as follows. We have
    $KO_{1}({\mathbb{R}})=KO_{0}(S{\mathbb{R}})$ $=\ker(KO_{0}(S{\mathbb{R}}^{+})\rightarrow KO_{0}({\mathbb{R}}))$
    $=\ker((\mathrm{ev}_{1})_{*}\colon KO_{0}(C^{\mathbb{R}}(S^{1}))\rightarrow KO_{0}({\mathbb{R}}))$
    The non-trivial element $\eta$ is represented by
    $\eta=\left[p\right]-\left[\left(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\right)\right]$
    where the projection $p\in M_{2}(C^{\mathbb{R}}(S^{1}))$ is defined by
    $p(x,y)=\tfrac{1}{2}\begin{pmatrix}1+x&y\\ y&1-x\end{pmatrix}$
    for $(x,y)\in S^{1}$. We check that $\mathrm{ev}_{1}(p)=\left(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\right)$, so that $\eta\in\ker(e_{1})_{*}$. Of course, the projection $p$ corresponds to the Möbius vector bundle over the circle. See Theorem 4.2 of Boersema [19]
  3. (3)
    We could also do the same thing for the non-trivial element of
    $KO_{2}({\mathbb{R}})=KO_{0}(S^{2}{\mathbb{R}}).$
    This time the corresponding projection defined on $(x,y,z)\in S^{2}$ is
    $p(x,y,z)=\tfrac{1}{2}\begin{pmatrix}1+x&0&y&z\\ 0&1+x&-z&y\\ y&-z&1-x&0\\ z&y&0&1-x\end{pmatrix}$
    See Theorem 4.4 of [19].
@@
    Now $SO_{2}\cong S^{1}$ and so $\pi_{1}(SO_{2})\cong{\mathbb{Z}}$, but $SO_{3}\cong RP^{3}$ as a topological space (where $RP^{3}$ is real projective space) and hence $\pi_{1}(SO_{3})={\mathbb{Z}}_{2}$. For $n>2$, any map $S^{1}\to SO_{n}$ is homotopic383838by the Cellular Approximation Theorem, cf. G. Whitehead [86] p. 77 and see §9 ahead. to a map taking values in $SO_{3}$ and hence $\pi_{1}(SO_{n})={\mathbb{Z}}_{2}$ for $n>2$, confirming that
@@ 7. External Products, the Künneth Theorem, and Internal Products
    $\mu_{O}\colon KO_{m}(A)\otimes KO_{n}(B)\rightarrow KO_{m+n}(A\otimes_{\mathbb{R}}B).$
@@
    $\mu_{U}\colon K_{m}(A)\otimes K_{n}(B)\rightarrow K_{m+n}(A\otimes_{\mathbb{C}}B).$
@@
Künneth theorems deal in general with determining $h_{*}(A\otimes B)$ in terms of $h_{*}(A)$ and $h_{*}(B)$. One has to put some constraints on the algebras in order for the theorem to hold. The full Künneth Theorem for $K$-theory involves some serious homological algebra. For $K_{*}(A)$ it was established by Schochet [73] (1984) where the bootstrap category was first introduced, and for $KO_{*}(A)$ it is due to Boersema [15] (2002), utilizing united $K$-theory. The following results are special cases of these general theorems. Anticipating events, we also state the corresponding results for the $K$-theory of topological spaces. See Section 11 for $K^{*}(X)$ and $K^{*}(X)$.
@@
    Suppose that $A$ and $B$ are complex $C^{*}$-algebras, $A$ is in the bootstrap category (which includes all commutative complex $C^{*}$-algebras), and $K_{*}(B)$ is torsion-free. Then the natural external product map induces an isomorphism of graded groups  [^40] If $G_{*}$ and $H_{*}$ are graded abelian groups then $(G_{*}\otimes H_{*})_{5}$ (for example) is understood to be the direct sum of the groups $G_{j}\otimes H_{5-j}$. This makes sense as long as only a finite number of these groups are non-zero. Bott periodicity makes things more complicated. So let us stipulate that whenever we are tensoring $K$-theory groups, that $\in\{0,1\}$ in the complex case and $\in\{0,1,\dots 6,7\}$ in the real case.
@@
    Suppose that $A$ and $B$ are real $C^{*}$-algebras, $A$ is in the pre-bootstrap category (this includes all commutative real $C^{*}$-algebras $A$), and $B$ has the property that $K_{*}(B_{\mathbb{C}})$ is torsion-free. Then the natural external product map induces an isomorphism of graded $KO_{*}({\mathbb{R}})$-modules [^41] We are casually assuming that $KO_{*}({\mathbb{R}})$ is a graded commutative ring that acts on $KO_{*}(A)$ for any real $C^{*}$-algebra $A$. We will establish this rigorously in Corollary 7.7. Actually the analogous statement in the complex case is also true, but since $K_{*}({\mathbb{C}}))$ is just $\mathbb{Z}$} and 0 in degrees 0 and $1$ respectively, the action has little significance here.
@@
    This includes cases where $B=C^{{\mathbb{R}}}(iS^{1})$, $C_{o}^{{\mathbb{R}}}(i{\bf{R}})$, $\real(S^{n})$. In particular, if $X$ and $Y$ are compact spaces and $Y$ has the homotopy type of a sphere or a product of spheres424242or a finite or countable disjoint union of such spaces, then there is a natural isomorphism
@@
From Theorem 6.18 we have
$KO_{n}({\mathbb{R}})\otimes{\mathbb{Q}}=\begin{cases}{\mathbb{Q}}&n=0,\,4\\ 0&\text{otherwise.}\end{cases}$
It turns out generally that $KO_{*}(A)\otimes{\mathbb{Z}}[\tfrac{1}{2}]$ and $KO_{*}(A)\otimes{\mathbb{Q}}$ are periodic with period $4$ for any real $C^{*}$-algebra. We will prove this in Section 24.
@@
Since ${\mathbb{H}}\otimes{\mathbb{H}}\cong M_{4}({\mathbb{R}})$ as remarked in §1, this simplifies to
@@
Corollary 7.11.
For any real $C^{*}$-algebra $A$, we have
$KH_{*}(A)\cong KO_{*+4}(A).$
Proof.
For any graded abelian group $M$, let $(\Sigma^{k}M)_{i}=M_{i+k}$. Then
$KH_{*}(A)$ $=KO_{*}(A\otimes{\mathbb{H}})$
$=KO_{*}(A)\otimes_{KO_{*}({\mathbb{R}})}KO_{*}({\mathbb{H}})$ by the Künneth Formula
$=KO_{*}(A)\otimes_{KO_{*}({\mathbb{R}})}\Sigma^{4}KO_{*}({\mathbb{R}})$ by Theorem 6.18
$=\Sigma^{4}(KO_{*}(A)\otimes_{KO_{*}({\mathbb{R}})}KO_{*}({\mathbb{R}}))$
$=\Sigma^{4}KO_{*}(A).$